Identify the slope and y- intercept of the line with the equation.
2x-3y=0

Answer by mananth(16946)

2x-3y=0
2x=3y
y=2x/3
y/x= 2/3
slope = 2/3
Here b = 0
in the equation y=2x/3 when comparing with y= mx+b
the line passes through the origin. No intercept